OpenStudy (anonymous):

Solve the proportion 2 times x all over 9 times x minus 8 equals 2 over x.. What is the value of x?

OpenStudy (anonymous):

Like this?\[\frac{ 2x }{ 9(x-8) } = \frac{ 2 }{ x }\]

OpenStudy (anonymous):

Or this?\[\frac{ 2x }{ 9x-8 } = \frac{ 2 }{ x}\]

OpenStudy (anonymous):

The first way has complex, imaginary roots, whereas the second way will produce two discrete, real roots.